### v2

# There is no definable Grauert Direct Image Theorem

Abstract.
We prove the claim in the title by showing that a definable Grauert Direct Image Theorem in o-minimal geometry would imply a weak representability-like property of the definable Picard functor. However, this weak representability cannot hold because of the definable Chow Theorem of Peterzil and Starchenko.
The second author was supported by the SFB 1085 Higher Invariants, Universität Regensburg
1. Introduction
Definable complex analytic geometry was initiated by Peterzil and Starchenko, see for example [PS09], who obtained fundamental results such as a definable analytic Noether normalization theorem and the famous definable Chow Theorem.
This foundational work on definable complex analytic geometry was continued by Bakker, Brunebarbe and Tsimerman in [BBT23, Sec. 2]. They applied it to prove Griffiths’ conjecture. In doing so, they define the notion of a definable complex analytic space. They then show that many of the basic results of complex analytic geometry hold in this context, in particular they prove the definable Oka Coherence Theorem [BBT23, Thm. 2.21] and the definable Grauert Direct Image Theorem for finite morphisms [BBT23, Prop. 2.52].
It was anticipated by experts that a general definable Grauert Direct Image Theorem might hold for arbitrary projective maps between algebraic varieties. We had hoped to be able to apply such a theorem in order to study moduli spaces of definable analytic vector bundles on compact Riemann surfaces and similar moduli spaces along the lines of [Sim94, p.112]. Unfortunately, it turns out that there can be no definable Grauert Direct Image Theorem for non-finite morphisms in any reasonable generality. This is the topic of this note.
Fix any o-minimal structure which expands $\mathbb{R}_{\rm an}$.
Theorem 1.1.
There exists a smooth projective definable morphism $f\colon X\to S$ of definable complex analytic spaces and a definable coherent sheaf $\mathcal{F}$ on $X$ such that $f_{*}\mathcal{F}$ is not definable coherent. More precisely, for $S=\mathbb{A}^{1}_{\mathbb{C}}\setminus\{0\}$, $X=Y\times S$ and $f$ the projection onto $S$, where $Y$ is a smooth projective algebraic variety with $H_{1}(Y,\mathbb{Q})\neq 0$, there always exists a semi-algebraically definable analytic line bundle $\mathcal{L}$ on $X$ such that the sheaves $f_{*}\mathcal{L}(n)$ for $n>0$ cannot all be definably coherent with respect to our fixed o-minimal structure.
In the theorem $f_{*}$ denotes the pushforward of sheaves on the definable sites relative to the fixed o-minimal structure.
Remark 1.2.
It seems possible that Theorem 1.1 holds true for all o-minimal structures. The only place in the proof where we use $\mathbb{R}_{\rm an}$ is Lemma 2.3.
The idea to prove Theorem 1.1 is to assume that the definable Grauert Direct Image Theorem is true and then to deduce a contradiction along the following lines. Consider a projective flat morphism $f\colon X\to S$ with geometrically integral fibers where $X$ and $S$ are complex algebraic varieties. For a definable line bundle $\mathcal{L}$ on $X$, the induced complex analytic map $\sigma_{\mathcal{F}}\colon S\to\mathrm{Pic}(X/S)$ to the Picard scheme would be definable if a definable Grauert Direct Image Theorem would hold, see Proposition 2.1.
However, it is not difficult to find such a geometric setting where $S$ is a rational variety, where $X=Y\times S$ with $Y$ a smooth projective algebraic variety and where $\mathcal{L}$ is the line bundle associated to a non-trivial family of rank one local systems on $Y$ parametrized by $S$, see Subsection 2.2. Then
$\sigma_{\mathcal{F}}\colon S\to\mathrm{Pic}(Y)\times S\cong\mathrm{Pic}(X/S)$
gives rise to a non-constant definable complex analytic morphism from $S$ to the abelian variety $\mathrm{Pic}^{0}(Y)$. By the definable Chow Theorem [PS09, Cor. 4.5] this morphism is algebraic, but there is no non-constant algebraic morphism from a rational variety $Z$ to an abelian variety $A$.
Acknowledgement. We thank Ben Bakker and Jacob Tsimerman for friendly comments on our construction and Bruno Klingler for a discussion.
2. The Construction
This section provides a detailed account of the construction outlined in the discussion following Theorem 1.1.
2.1. Definability of sections of the Picard scheme
In the following we fix an o-minimal structure which expands $\mathbb{R}_{\rm an}$. Throughout the subsequent discussion, the Picard scheme may be replaced, in greater generality, with the coarse moduli scheme of semi-stable coherent sheaves.
For a definable complex analytic space $U$ we denote by $\mathcal{O}_{U}$ the sheaf of definable complex analytic functions. A definable line bundle $\mathcal{L}$ is a sheaf of $\mathcal{O}_{U}$-modules such that there is a finite covering $U=\cup_{i\in I}U_{i}$ by definable open subsets $U_{i}\subset U$ with $\mathcal{L}|_{U_{i}}\cong\mathcal{O}_{U_{i}}$ for all $i\in I$. We denote by $\mathcal{L}^{\rm an}$ the underlying complex analytic line bundle.
Consider a flat projective morphism with geometrically integral fibers $f\colon X\to S$ and with $X$ and $S$ algebraic varieties. Let $\mathrm{Pic}(X/S)$ be Grothendieck’s Picard scheme [Gro95]. It is well-known that for any analytic open subset $U\subset S$ and an analytic line bundle $\mathcal{L}$ on $X_{U}=f^{-1}(U)$ the induced map on $\mathbb{C}$}-points $\sigma_{\mathcal{L}}\colon U\to\mathrm{Pic}(X/S)|_{U}$ is holomorphic, see [Sim94, Sec. 5].
Proposition 2.1.
Let $\mathcal{L}$ be a definable line bundle on $X_{U}$. If for all integers $n>0$ the sheaf $f_{*}\mathcal{L}(n)$ on $U$ is definable coherent, then the section $\sigma_{\mathcal{L}}$ is definable.
Note that the image of $\sigma_{\mathcal{L}}$ is contained in an open and closed subscheme of $\mathrm{Pic}(X/S)$ which is of finite type over $S$, given by fixing finitely many Hilbert polynomials, so the notion of definability makes sense here. The idea of the proof of Proposition 2.1 is simply to copy [Sim94, Sec. 5] and add the word definable when needed. We are going to present a concise version of Simpson’s construction.
Proof.
Assume that the sheaves $f_{*}\mathcal{L}(n)$ are definable coherent for all $n>0$. The canonical map
$(f_{*}\mathcal{L}(n))^{\rm an}\xrightarrow{\sim}f_{*}\mathcal{L}^{\rm an}(n)$
is an isomorphism by Lemma 2.3. By Grauert’s Base Change Theorem [GR84, Sec. 10.5.5] and Lemma 2.2 there exists an integer $n$ such that
$f^{*}f_{*}\mathcal{L}^{\rm an}(n)\to\mathcal{L}^{\rm an}(n)$
is surjective. As $f_{*}\mathcal{L}(n)$ is definable coherent, after shrinking $U$ there is a surjection
$\oplus_{\{1,\ldots,b\}}\mathcal{O}_{X_{U}}\to f_{*}\mathcal{L}(n)$
which thus induces a surjection
$\oplus_{\{1,\ldots,b\}}\mathcal{O}_{X_{U}}(-n)\to\mathcal{L}.$
Let $\mathcal{K}$ be the kernel of this map. Note that $\mathcal{K}$ is definable coherent [BBT23, Lem. 2.14]. Again by Grauert’s Base Change Theorem and Lemma 2.2 there exists an integer $m$ such that
$R^{i}f_{*}\mathcal{K}^{\rm an}(m)=R^{i}f_{*}\mathcal{O}^{\rm an}_{X_{U}}(m-n)=R^{i}f_{*}\mathcal{L}^{\rm an}(m)=0$
for all $i>0$. This implies that the analytification of
$0\to f_{*}\mathcal{K}(m)\to\oplus_{\{1,\ldots,b\}}f_{*}\mathcal{O}_{X_{U}}(m-n)\to f_{*}\mathcal{L}(m)\to 0$
is an exact sequence of locally free $\mathcal{O}_{U}^{\rm an}$-sheaves, so this is a short exact sequence of definable locally free coherent sheaves itself by [BBT23, Thm. 2.27] and Lemma 2.5.
Let $r$ be the rank of $f_{*}\mathcal{L}(m)$. This short exact sequence induces a definable analytic map to a Grassmannian classifying codimension $r$ locally free subsheaves
$\tilde{\sigma}_{\mathcal{L}}\colon U\to\mathrm{Grass}_{U}(\oplus_{\{1,\ldots,b\}}f_{*}\mathcal{O}_{X_{U}}(m-n),r).$
By the general construction of moduli spaces of coherent algebraic sheaves, there is an open subscheme $\mathrm{Quot}^{\prime}$ of finite type over $S$ of Grothendieck’s Quot scheme classifying coherent quotients of $\oplus_{\{1,\ldots,b\}}\mathcal{O}_{X}(-n)$. Here $\mathrm{Quot}^{\prime}$ classifies line bundle quotients with the same Hilbert polynomial as the line bundles $\mathcal{L}_{s}$ (assume $S$ is connected for simplicity).
For suitable $m$ the above construction provides a locally closed immersion
$\mathrm{Quot}^{\prime}\hookrightarrow\mathrm{Grass}_{S}(\oplus_{\{1,\ldots,b\}}f_{*}\mathcal{O}_{X}(m-n),r).$
Consider the canonical scheme morphism $q\colon\mathrm{Quot}^{\prime}\to\mathrm{Pic}(X/S)$. As $\sigma_{\mathcal{L}}=q\circ\tilde{\sigma}_{\mathcal{L}}$ is the composition of definable maps, we finally conclude that $\sigma_{\mathcal{L}}$ is definable. ∎
The following lemma is a standard variant of Serre’s vanishing theorem. It is a consequence of a combination of Thm. III.5.2 and of Thm. III.12.11(b) in [Har77].
Lemma 2.2.
For a bounded family of line bundles $\mathcal{L}^{(\lambda)}_{s_{\lambda}}$ on fibers $X_{s_{\lambda}}$, $\lambda\in\Lambda$, there exists an $n>0$ such that $\mathcal{L}^{(\lambda)}_{s_{\lambda}}(n)$ is generated by global sections and such that $H^{i}(X_{s_{\lambda}},\mathcal{L}^{(\lambda)}_{s_{\lambda}}(n))$ vanishes for all $i>0$.
The proof of the following Lemma 2.3 is the only place where we use that our o-minimal structure expands $\mathbb{R}_{\rm an}$. It seems possible that the lemma is true more generally.
Lemma 2.3.
Let $\mathcal{F}$ be a definable coherent sheaf on $X$. Then the canonical map
$\psi\colon(f_{*}\mathcal{F})^{\rm an}\xrightarrow{\sim}f_{*}(\mathcal{F}^{\rm an})$
is an isomorphism.
Proof.
It suffices to show that for any $\mathbb{C}$}-point $s\in S$ the induced map $\psi_{s}$ on stalks is an isomorphism. Its injectivity is clear. For surjectivity consider $\theta\in(f_{*}\mathcal{F}^{\rm an})_{s}$. Say $\theta$ is induced by $\tilde{\theta}\in\mathcal{F}^{\rm an}(X_{V})$, where $V\subset S$ is a definable analytic open subset containing $s$. Let $V^{\prime}\Subset V$ be a relatively compact definable open subset containing $s$. Then by Claim 2.4 $\tilde{\theta}|_{X_{V^{\prime}}}$ is induced by an element in $\mathcal{F}(X_{V^{\prime}})$, which gives rise to a preimage of $\theta$ in $(f_{*}\mathcal{F})_{s}=(f_{*}\mathcal{F})^{\rm an}_{s}$.
Claim 2.4.
For a definable coherent sheaf $\mathcal{G}$ on a definable analytic space $V$ and for $V^{\prime}\Subset V$ a relatively compact definable open subset the image of the restriction map $\mathcal{G}^{\rm an}(V)\to\mathcal{G}^{\rm an}(V^{\prime})$ has values in $\mathcal{G}(V^{\prime})\subset\mathcal{G}^{\rm an}(V^{\prime})$.
The conclusion of the claim is local, so we can assume without loss of generality that there is a surjection $\psi\colon\mathcal{O}_{V}^{b}\to\mathcal{G}$ and that we have to show the claim for an element of the form $\psi(\zeta)$. In other words we are reduced to showing the claim for $\mathcal{G}$ the structure sheaf and for $V\subset\mathbb{C}^{n}$ a basic definable complex analytic space [BBT23, Def. 2.19]. In this case the claim is clear by the definition of $\mathbb{R}_{\rm an}$. ∎
Lemma 2.5.
If $\mathcal{F}$ is a definable coherent sheaf on a definable analytic space $V$ such that $\mathcal{F}^{\rm an}$ is locally free, then $\mathcal{F}$ is locally free.
Proof.
We find a finite definable open covering $V=\cup_{i\in I}V_{i}$ together with surjective morphisms $\psi_{i}\colon\mathcal{\oplus}_{\{1,\ldots,b\}}\mathcal{O}_{V_{i}}\to\mathcal{F}|_{V_{i}}$. We can assume that $\mathcal{F}|_{V_{i}}^{\rm an}$ is locally constant of rank $r_{i}$. For a subset $W\subset\{1,\ldots,b\}$ of cardinality $r_{i}$, let $V_{i,W}\subset V_{i}$ be the open subset over which $\psi_{i}|_{\oplus_{W}\mathcal{O}_{V_{i}}}$ is an isomorphism. Then $V_{i,W}$ is the complement of the union of the supports of $\mathrm{Ker}(\psi_{i})$ and ${\rm Coker}(\psi_{i})$. By [BBT23, Lem.2.14] it is thus definable. Then $V=\cup_{i\in I,W}V_{i,W}$ is a finite definable open covering such that $\mathcal{F}|_{V_{i,W}}$ is free. ∎
2.2. The character variety and definable local systems of rank one
Let $Y$ be a complex projective variety. We fix an arbitrary o-minimal structure. Its character variety
${\rm Char}(Y)={\rm Hom}(\pi^{\rm ab}_{1}(Y),{\mathbb{G}}_{m})$
is a diagonalizable linear algebraic group which classifies complex local systems of rank one on $Y$.
For a definable analytic space $U$ we write
[…]
Here $X_{U}=Y\times U$ and $f\colon X_{U}\to U$ is the projection. Note that $\mathrm{M}(U)$ is an ad hoc definable variant of the framed moduli of local systems [Sim94, Lem. 7.3], which is sufficient for our purpose.
Proposition 2.6.
The canonical map which associates to an element of $M(U)$ the family of complex representation of $\pi_{1}^{\rm ab}(Y)$ parametrized by $U$
$\Psi\colon\mathrm{M}(U)\xrightarrow{\sim}\mathrm{Hom}(U,{\rm Char}(Y))$
is bijective. Here the codomain consists of all definable analytic maps.
Proof.
Let $\mathcal{V}=(V_{i})_{i\in I}$ be a finite definable open covering of $Y$. By [vdD98, Ch. 8] we can find such a covering with all $V_{i}$ contractible. Let $\check{Z}^{1}=\check{Z}^{1}(\mathcal{V},\mathbb{C}^{\times})$ be the group of Čech $1$-cocycles, which has the structure of a complex diagonalizable linear algebraic group. Similarly, $\check{C}^{0}=\check{C}^{0}(\mathcal{V},\mathbb{C}^{\times})$ has the structure of an algebraic torus, and we obtain an exact sequence of abelian linear algebraic groups
(2.1) $\check{C}^{0}\xrightarrow{d}\check{Z}^{1}\to{\rm Char}(Y)\to 1$
where the map on the right associates to a local system its monodromy representation. The canonical map
(2.2) $\mathrm{Hom}(U,\check{Z}^{1})/\mathrm{Hom}(U,\check{C}^{0})\to\mathrm{M}(U)$
which associates a locally free $f^{-1}\mathcal{O}_{U}$-sheaf of rank one to a Čech cocycle in
$\mathrm{Hom}(U,\check{Z}^{1}(\mathcal{V},\mathbb{C}^{\times}))=\check{Z}^{1}((V_{i}\times U)_{i\in I},f^{-1}\mathcal{O}_{U}^{\times})$
is bijective in view of the standard correspondence between the first Čech cohomology and the set of isomorphism classes of locally free sheaves. Surjectivity in (2.2) holds as any locally free definable $f^{-1}\mathcal{O}_{U}$-sheaf on $X_{U}$ is trivial on $V_{i}\times U$ for all $i\in I$ as $V_{i}$ is simply connected.
Because the exact sequence (2.1) splits as a sequence of algebraic tori if we pass to the connected components of the neutral element, there are algebraic sections ${\rm Char}(Y)\to\check{Z}^{1}$ and $\operatorname{{\rm im}}(d)\to\check{C}^{0}$ of the maps in (2.1). Via these sections we deduce that the sequence of groups of definable morphisms
$\mathrm{Hom}(U,\check{C}^{0})\xrightarrow{d}\mathrm{Hom}(U,\check{Z}^{1})\to\mathrm{Hom}(U,{\rm Char}(Y))\to 1$
is exact. Combined with (2.2) this finishes the proof. ∎
2.3. Proof of Theorem 1.1
As we assume that $H_{1}(Y,\mathbb{Z})\cong\pi_{1}^{\rm ab}(Y)$ is not a torsion group, it follows that $\mathrm{Char}(Y)$ has positive dimension. So we find a monomorphism $\mathbb{G}_{m}\to\mathrm{Char}(Y)$ of algebraic groups which gives rise by Proposition 2.6 to a definable locally free $f^{-1}\mathcal{O}_{\mathbb{G}_{m}}$ sheaf $\mathbb{L}$ of rank one on $X=Y\times{\mathbb{G}_{m}}$. Consider the associated definable line bundle $\mathcal{L}=\mathbb{L}\otimes_{f^{-1}\mathcal{O}_{\mathbb{G}_{m}}}\mathcal{O}_{X}$ and the associated analytic map to the Picard scheme $\sigma_{\mathcal{L}}\colon\mathbb{G}_{m}\to\mathrm{Pic}(Y)$.
Claim 2.7.
The map $\sigma_{\mathcal{L}}$ is non-constant.
Proof.
For $s\in\mathbb{G}_{m}$ with $|s|=1$ the local system $\mathbb{L}_{s}$ on $Y$ is unitary. Non-isomorphic unitary local systems of rank one give rise to non-isomorphic line bundles by Hodge theory. So the images by $\sigma_{\mathcal{L}}$ of these infinitely many unitary $s$ are pairwise different, i.e. $\sigma_{\mathcal{L}}$ is not constant. ∎
Finally, if the definable sheaves $f_{*}\mathcal{L}(n)$ were coherent for all $n>0$ then $\sigma_{\mathcal{L}}$ would be definable by Proposition 2.1. But then $\sigma_{\mathcal{L}}$ would be algebraic by the definable Chow theorem [PS09, Cor. 4.5]. However, algebraic maps from a rational variety to an abelian variety are constant, see for example [Mil86, Cor. 3.9], so $\sigma_{\mathcal{L}}$ is constant contradicting Claim 2.7.